The following data represent the age​ (in weeks) at which babies first crawl based on a survey of 12 mothers. The data are normally distributed and sequals9.930 weeks. Construct and interpret a 99​% confidence interval for the population standard deviation of the age​ (in weeks) at which babies first crawl.

55 31 43 35 39 26 48 37 53 27 39 27
Answer #1

Solution:

Confidence interval for population standard deviation is given as below:

Sqrt[(n – 1)*S2 / χ2α/2, n– 1 ] < σ < sqrt[(n – 1)*S2 / χ21 -α/2, n– 1 ]

From given data, we have

Confidence level = 99%

Sample size = n = 12

Degrees of freedom = n – 1 = 11

Sample standard deviation = S = 9.93

χ2α/2, n – 1 = 26.7568

χ21 -α/2, n– 1 = 2.6032

(By using chi square table)

Sqrt[(12 – 1)* 9.93^2 / 26.7568] < σ < sqrt[(12 – 1)* 9.93^2 / 2.6032]

Sqrt[40.5374] < σ < sqrt[416.6583]

6.3669 < σ < 20.4122

Lower limit = 6.3669

Upper limit = 20.4122